OpenStudy (anonymous):

please help! Calculus is going to be the death of me. find the linear approximation of the function at a. f(x)=sinx a=pi/6 I am suppose to use square root 3/2, 1/2 instead of pi/6

OpenStudy (anonymous):

is your question to find the linear approximation for f(1/2) ???

OpenStudy (anonymous):

if it is, then you can use the equation of the tangent line at x=pi/6 to approximate f(1/2)=sin(1/2)

OpenStudy (anonymous):

well i know the formula is L(x)=f(a)+f'(a)(x-a) do i just plug in 1/2, and why do I use 1/2 instead of square root3/2

OpenStudy (anonymous):

f(a)=sin(pi/6) = 1/2 f'(a)=cos(pi/6) = \(\sqrt3/2 \) a = pi/6 L(x) is the equation of the tangent line at x=pi/6 (in slope-intercept form)

OpenStudy (anonymous):

i see... the tangent line is the linear approximation of f(x)=sinx at x=pi/6......

OpenStudy (anonymous):

that makes a lot of sense, I was making it a lot more complicated than it needed to be. Thanks for helping me understand.

OpenStudy (anonymous):

yw.... :)